OpenStudy (anonymous):

2x-3(4-x)=3+4(x-1) Solve for X

OpenStudy (anonymous):

Open the parentheses using the distribution law: \(x(y+z)=xy+xz\). This tells you how to handle expressions which combine products and sums. That's actually the challenging part here, expanding the parts \(-3(4-x)\) and \(3(x-1)\). Then transpose all x to one side of the equations and constant numbers to another side.

OpenStudy (anonymous):

So then the equation becomes 8x-12-2xsquared=3x-3+4x-4 ?

OpenStudy (anonymous):

There won't be anything squared. \(2x-3(4-x)=3+4(x-1)\) \(2x-12+3x=3+4x-4\) \(5x-12=-1+4x\) \(5x-4x=-1+12\) \(x=11\)